OpenStudy (anonymous):

Say you have a right triangle with side a=21 and side b=220. what is the length of side c?

OpenStudy (jdoe0001):

use the pythagorean theorem \(\bf c^2 = a^2+b^2 \implies c = \sqrt{a^2+b^2}\)